# EgoEMS: A High-Fidelity Multimodal Egocentric Dataset for Cognitive Assistance in Emergency Medical Services

### v1

@@ 2 Background and Related Work
Despite these advances, existing datasets are confined to routine tasks, controlled surgical scenes, or patient-centric contexts, and none capture the procedural structure, high‑stakes environment, and coordinated multiperson nature of end‑to‑end EMS workflows (see Table 1). Curation of such datasets is particularly challenging due to privacy concerns and high annotation costs in medical settings, resulting in much smaller datasets. No prior dataset offers multimodal, multiperson egocentric recordings of simulated emergencies with certified responders and the ground truth necessary for modeling decision‑making and cognitive assistance in critical care.
@@
In this paper, we take the first step towards benchmarking SOTA deep learning models for activity recognition using multimodal data for the EMS domain. We select strong, representative baselines that cover complementary approaches, including supervised models based on transformer architectures gberta_2021_ICML; weerasinghe2024multimodal capable of modeling long temporal dependencies and multimodal fusion, a convolutional TSN model wang2016temporal as a widely used CNN baseline, and zero-shot vision–language models such as Qwen-2.5 bai2025qwen2 and VideoLLaMA-3.3 zhang2025videollama to assess the potential of large pretrained models. For audio-only recognition, we also include a custom zero-shot pipeline using WhisperTimestamped lintoai2023whispertimestamped and GPT-4o achiam2023gpt.
@@ 3.1 EMS Taxonomy
Figure 2: Methodology for creating the EgoEMS dataset including the EMS taxonomy.
@@ 3.2 Simulation Experiments
We conducted 233 simulated EMS trials spanning approximately 20 hours, encompassing a range of cardiac and stroke related emergency scenarios. The simulations were divided into two primary types: high-fidelity scenarios, where procedures were performed on human actors portraying patients, and cardiac arrest scenarios, which used manikins due to the nature of CPR intervention. High-fidelity simulations featured patients with diverse medical histories and demographics to promote data diversity and generalizability. Each simulation captured end-to-end EMS procedures performed by a team of 2–3 responders, performing the critical interventions shown in Table 2. In addition, several scenarios were designed to reflect complex, realistic cases in which the initial chief complaint and presenting neurological symptoms mimicked a stroke but were ultimately attributable to alternative causes (e.g., hypoglycemia). These cases required responders to accurately assess, differentiate, and respond using appropriate protocol guided decision making. In cardiac arrest trials, EMS responders typically operated in pairs designated as primary and secondary responders carrying out critical interventions such as CPR, ventilation, and defibrillation on a manikin. In contrast, subjects from the general public conducted the trials individually and performed only CPR intervention on a manikin due to lack of medical training. To further enhance realism, additional volunteers served as bystanders, particularly in scenarios where the patient was unresponsive, contributing information such as patient history and situational context.
@@ 3.2.2 Privacy and Ethics
We obtained IRB approval prior to data collection, adhering to ethical standards for human subject research. We further discuss privacy and ethics challenges in real-world settings beyond simulation data in Appendix B along with details on the IRB and the data de-identification process.
@@ 3.4 Annotations
We employed manual and semi‑automatic approaches, with some annotations purely manual and others using zero‑shot models with manual verification. Table 2 shows a summary of annotations in the dataset. See Appendix B for details.
Figure 3: Top-10 keystep distribution with average duration. See Appendix A for a complete distribution.
Fine-grained keystep annotation. EgoEMS is manually annotated for 67 keysteps belonging to 9 interventions (see Appendix B and Figure 3). The dataset captures multiperson activity typical of real-world EMS procedures, where responders often perform concurrent actions during the same time interval (e.g., AED activation during chest compressions; see Appendix A). While the dataset includes multiperson annotations, our benchmarks focus exclusively on the primary responder and their egocentric viewpoint. We do not leverage secondary responder annotations and multiview data for model training or evaluation, but include them to support future research on multiperson activity understanding.
Timestamped speaker diarization and transcription. We designed an automated LLM based pipeline to generate speaker-diarized timestamped transcripts for each trial’s audio, which were then manually verified (see Appendix A). The general public participants struggled to narrate while performing CPR due to lack of advanced training, so they were instructed to focus solely on performing the task.
Bounding box and segmentation mask annotations. We generated bounding box and segmentation mask annotations for medical objects involved in cardiac arrest interventions (see Figure 1) using a semi-automatic pipeline. The nature of these interventions involves frequent and sustained interactions with critical medical equipment, making object localization particularly relevant in this context, serving as a promising candidate for improving the activity recognition capability of an ICA. This pipeline leverages a SOTA object detection model fine-tuned based on our EMS taxonomy combined with a zero-shot segmentation method. Our manual verification of 10% of the bounding box annotations against human annotations shows this method saves over 60 hours of annotation time at a slight loss of precision.
@@
Figure 4: Ground-truth CPR compression rates for participants with varying skill levels. Each subject’s reported rate is averaged across all their trials.
Figure 5: Overview of benchmark tasks for keystep classification, segmentation, and CPR quality estimation using multimodal data from EgoEMS. Example baseline models and representative results are shown to highlight the diverse evaluation scenarios enabled by the dataset.
@@ 4 EgoEMS Benchmarks
We present three benchmark tasks designed to evaluate the core capabilities of an ICA for EMS: (1) Keystep Classification and (2) Keystep Segmentation, which together form the broader task of keystep recognition, a foundational capability for guiding responders through complex protocols while monitoring their actions and (3) Action Quality Estimation, which enables continuous feedback to improve execution quality. As illustrated in Figure 5, these tasks leverage synchronized multimodal data including egocentric video, audio, and smartwatch IMU to support real-time inference of procedural context and responder performance. Below, we provide an overview of each benchmark; details and baseline results are included in Appendix D.
@@ 5 Conclusion
EgoEMS is the first egocentric, multimodal, multiperson dataset dedicated to EMS, created to drive the development of AI systems that can function as virtual partners to first responders in the field. Developed in close collaboration with EMS professionals and aligned with national standards, it captures high‑fidelity simulations of critical EMS procedures. EgoEMS provides over 20 hours of synchronized multimodal data, including 9 interventions and 67 keysteps across 233 trials, performed by 62 participants from multiple EMS agencies and the general public. The dataset’s comprehensive taxonomy and fine-grained annotations are complemented by benchmark tasks that showcase the potential of cognitive assistants and support exploration of multimodal fusion. Furthermore, we provide open-source resources that enable reproducibility and future extensions. EgoEMS establishes a new benchmark for multimodal AI research and lays a robust foundation for next-generation EMS technologies that enhance responder performance, reduce cognitive burden, and improve patient outcomes.
@@ Acknowledgments
This work was supported in part by the award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST) and a research grant from the Commonwealth Cyber Initiative (CCI). We are grateful for the support and participation of several volunteer EMS first responders at rescue squads and fire agencies in the Charlottesville and Virginia Beach areas, specifically North Garden Volunteer Fire Company, Western Albemarle Rescue Squad, Charlottesville Albemarle Rescue Squad and Ocean Park Volunteer Rescue Squad, as well as George Stephens, James Fitz-Gerald, Margaret Sande, Jon Howard, and David Keeler, who helped us with organizing the simulation experiments.
@@ A.1 NEMSIS
Intervention Keystep Description
check_pulse Assess for carotid or radial pulse.
check_breathing Look, listen, and feel for breathing.
chest_compressions Perform compressions at correct rate/depth.
assess_patient Reassess patient condition.
Defibrillation request_aed Request AED early in resuscitation.
request_assistance Call for bystander or team assistance.
turn_on_aed Power on AED and prepare for use.
attach_defib_pads Attach pads to bare chest correctly.
clear_for_analysis Ensure no contact during rhythm analysis.
clear_for_shock Confirm everyone is clear before shock.
administer_shock_aed Deliver shock if advised by AED.
Ventilation open_airway Use airway technique (e.g., head-tilt).
place_bvm Position and seal BVM mask.
ventilate_patient Deliver ventilations appropriately.
suction_airway Use suction to clear airway.
inset_NPA Insert NPA to maintain airway.
@@ A.2 Keysteps
Following selecting the protocols and interventions, we consulted the National Registry of Emergency Medical Technicians (NREMT) to review the keysteps and guidelines necessary for accurately executing each intervention. To ensure the robustness of our taxonomy, we collaborate closely with EMS experts, who validate the defined keysteps for each procedure. Specifically, we reference the Emergency Medical Responder Psychomotor Examination criteria nremt_psychomotor_exams to inform the development of this keystep taxonomy, aligning it with established standards for practical skill assessment. Tables 4,6,5 present the taxonomy of EMS interventions and keysteps defined for our dataset.
@@ A.3 Scenarios
This structure cascades into specific interventions (e.g., Vital Sign Assessment, 12-lead ECG, Transport), each decomposed into granular keysteps (e.g., explain procedure, place V1 lead, interpret ECG). To support model generalization and realism, we designed diverse scenarios featuring patients with varying comorbidities and sociocultural backgrounds. Scenario briefs were provided to participants before each simulation to standardize objectives while allowing natural variation in responder-patient interactions. For example, in the context of cardiac arrest emergency, we created multiple scenarios (see Figure 8) featuring patients with varied backgrounds and comorbidities, thereby enhancing the diversity and ecological validity of the dataset. These scenario outlines were shared with participants prior to each simulation to guide the general flow of actions and conversations, while still allowing for responders to dynamically react to patient condition.
@@ A.4 Annotation
Figure 10: Video annotation tool from Oxford dutta2019vgg.
@@ Appendix D Benchmarks
Model (Architecture) Modality (Feature Extraction) Classification Metric
Precision Recall F1-score Top-1 Acc. Top-5 Acc.
MTRSAP (Transformer) weerasinghe2024multimodal Ego Video (Resnet50) 0.55 0.58 0.55 62.27 77.01
Ego Video (I3D) 0.56 0.57 0.56 60.41 72.50
Ego Video, IMU (Resnet50, Norm.) 0.55 0.58 0.56 62.19 76.14
IMU (Norm.) 0.42 0.45 0.42 48.34 67.57
Ego Video, Audio (Resnet50, Mel-Spec) 0.19 0.23 0.20 25.67 44.38
Ego Video, Audio (Resnet50, Wav2Vec2) 0.34 0.39 0.35 44.71 66.16
Audio (Wav2Vec2) 0.26 0.27 0.26 33.25 52.14
VideoLlama-3.3 (VLM)
zhang2025videollama Ego Video (End-to-End) 0.16 0.17 0.16 18.51 -
Qwen-2.5 (VLM)
bai2025qwen2 Ego Video (End-to-End) 0.33 0.35 0.34 38.30 -
GPT-4o (LLM) achiam2023gpt Audio (End-to-End) 0.17 0.19 0.18 23.89 -
MM-CDFSL (Transformer) Hatano2024MMCDFSL Ego Video (Few-Shot, MM Distilled) $\dagger$ - - - 67.70 -
@@ D.1.2 Multimodal Few-Shot Setting
Knowledge transfer and multimodal distillation show promising performance in low-data regimes; to assess cross-domain transfer, We evaluate cross-domain, multimodal few-shot learning with MM-CDFSL Hatano2024MMCDFSL. Following the authors, we use the distilled RGB student encoder on a large scale general purpose egocentric dataset (Ego4D grauman2022ego4d) (teachers: optical flow and hand pose) and perform episodic 5-way, 5-shot evaluation on our dataset. For each episode, a linear classifier (logistic regression) is fit on the support set and evaluated on the query set using RGB only; we also apply ensemble masked inference as in the original work. We report the mean top-1 accuracy across 600 episodes.

### v2

@@ 2 Background and Related Work
Despite these advances, none of the existing datasets capture the procedural structure, high‑stakes environment, and coordinated multiperson nature of end‑to‑end EMS workflows (see Table 1). Curation of such datasets is particularly challenging due to privacy concerns and high annotation costs in medical settings, resulting in much smaller datasets. No prior dataset offers multimodal, multiperson egocentric recordings of simulated emergencies with certified responders and the ground truth necessary for modeling decision‑making and cognitive assistance in critical care.
@@
In this paper, we take the first step towards benchmarking SOTA deep learning models for multimodal EMS activity recognition. We evaluate diverse representative models, covering supervised (unimodal and multimodal), cross-domain few-shot, and zero-shot (LLM) approaches. See EgoEMS Benchmarks Section and Appendix D for more details.
@@ 3.1 EMS Taxonomy
Figure 2: Methodology for creating the EMS taxonomy and EgoEMS simulation scenarios.
@@ 3.2 Simulation Experiments
We conducted 233 simulated EMS trials spanning approximately 20 hours, encompassing a range of cardiac and stroke related emergency scenarios. The simulations were divided into two primary types: (i) high-fidelity scenarios, where procedures were performed on human actors portraying patients, and (ii) cardiac arrest scenarios, which used manikins for performing CPR interventions. High-fidelity simulations were designed to emphasize realism in settings and workflows (e.g., ambulance transfers, in-vehicle recordings with equipment noise), human dynamics (e.g., interfering bystanders), and patient variability, incorporating diverse medical histories and demographics to enhance data diversity. Each simulation captured end-to-end EMS procedures performed by a team of 2–3 responders, performing the critical interventions shown in Table 2. In addition, several scenarios were designed to reflect complex, realistic cases in which the initial chief complaint and presenting neurological symptoms mimicked a stroke but were ultimately attributable to alternative causes (e.g., hypoglycemia). These cases required responders to accurately assess, differentiate, and respond using appropriate protocol guided decision making. In cardiac arrest trials, EMS responders typically operated in pairs designated as primary and secondary responders carrying out critical interventions such as CPR, ventilation, and defibrillation on a manikin. In contrast, subjects from the general public conducted the trials individually and performed only CPR intervention on a manikin due to lack of medical training. To further enhance realism, additional volunteers served as bystanders, particularly in scenarios where the patient was unresponsive, contributing information such as patient history and situational context.
@@ 3.2.2 Privacy and Ethics
We obtained IRB approval prior to data collection and adhered to ethical and privacy protection guidelines for human subjects research in all aspects of recruitment, experiments, and data management. All identifying information, including participants’ names, faces, ID cards, and vehicles’ license-plates were removed from video and audio before data release. More details on the IRB protocol, de-identification methods, and real-world ethical considerations are in Appendix B and the Ethical Statement.
@@ 3.4 Annotations
We employed manual and semi‑automatic approaches, with some annotations purely manual and others using zero‑shot models with manual verification. Table 2 shows a summary of annotations in the dataset. See Appendix A for details.
Figure 3: Top-10 keystep distribution with average duration.
Fine-grained keystep annotation. EgoEMS is manually annotated for 67 keysteps belonging to 9 interventions (see Appendix A.2 and Figure 3). The dataset captures multiperson activity typical of real-world EMS procedures, where responders perform concurrent actions during the same time interval (e.g., AED activation during chest compressions; see Appendix A). While the dataset includes multiperson annotations, our benchmarks focus on the primary responder and their egocentric viewpoint. We do not leverage secondary responder annotations and multiview data for model training or evaluation, but include them to support future research on multiperson activity understanding.
Timestamped speaker diarization and transcription. We designed an automated LLM based pipeline to generate speaker-diarized timestamped transcripts for each trial’s audio, which were then manually verified (see Appendix A.4.2). The general public participants struggled to narrate while performing CPR due to lack of advanced training, so they were instructed to focus solely on performing the task.
Bounding box and segmentation mask annotations. We generated bounding box and segmentation mask annotations for medical objects involved in cardiac arrest interventions (see Figure 1) using a semi-automatic pipeline. The nature of these interventions involves frequent and sustained interactions with critical medical equipment, making object localization particularly relevant in this context, serving as a promising candidate for improving the activity recognition capability of an ICA. This pipeline leverages a SOTA object detection model fine-tuned based on our EMS taxonomy combined with a zero-shot segmentation method. Our manual verification of 10% of the bounding box annotations against human annotations shows this method saves over 60 hours of annotation time at a slight loss of precision. We refer the reader to Appendix A.4.3 for more detailed analysis.
@@
Figure 4: Ground-truth average CPR compression rate per participant across three skill levels (Public, Member, EMT).
Figure 5: Overview of benchmark tasks for keystep classification, segmentation, and CPR quality estimation using EgoEMS data. Example baseline models and representative results highlight the diverse evaluation scenarios enabled by the dataset.
@@ 4 EgoEMS Benchmarks
We present three benchmark tasks designed to evaluate the core capabilities of an ICA for EMS: (1) Keystep Classification and (2) Keystep Segmentation, which together form the broader task of keystep recognition, a foundational capability for guiding responders through complex protocols while monitoring their actions and (3) Action Quality Estimation, which enables continuous feedback to improve execution quality. As illustrated in Figure 5, these tasks leverage synchronized multimodal data including egocentric video, audio, and smartwatch IMU to support real-time inference of procedural context and responder performance.
Below, we provide an overview of each benchmark task and results from a representative set of SOTA models. We select strong, representative baselines that cover complementary approaches, including transformer architectures gberta_2021_ICML; weerasinghe2024multimodal capable of modeling long temporal dependencies and multimodal fusion, a convolutional TSN model wang2016temporal as a widely used CNN baseline, few-shot cross-domain models such as MM-CDFSL Hatano2024MMCDFSL and zero-shot vision–language models such as Qwen-2.5 bai2025qwen2 and VideoLLaMA-3.3 zhang2025videollama to assess the potential of large pretrained models. For audio-only recognition, we also include a custom zero-shot pipeline using WhisperTimestamped lintoai2023whispertimestamped and GPT-4o achiam2023gpt. More detailed results and discussions are in Appendix D.
@@ 5 Conclusion
EgoEMS is the first end-to-end high-fidelity egocentric, multimodal, multiperson dataset, created to drive the development of AI systems that can support EMS first responders in the field. Developed in close collaboration with EMS professionals and aligned with national standards, it provides over 20 hours of synchronized multimodal data from IRB-approved simulated emergencies with real responders, including 9 interventions and 67 keysteps across 233 trials performed by 62 participants from multiple EMS agencies and the general public. EgoEMS’s low-cost and easily replicable data collection system, comprehensive taxonomy, fine-grained annotations, and open-source annotation and de-identification tools establish a foundation for developing and expanding realistic and high-fidelity medical datasets and multimodal cognitive assistants.
@@ Acknowledgments
This work was supported in part by the award 70NANB21H029 from the U.S. Department of Commerce, National Institute of Standards and Technology (NIST), and a research grant from the Commonwealth Cyber Initiative (CCI). We are grateful for the support and participation of several volunteer EMS first responders at rescue squads and fire agencies in the Charlottesville and Virginia Beach areas, specifically North Garden Volunteer Fire Company, Western Albemarle Rescue Squad, Charlottesville Albemarle Rescue Squad and Ocean Park Volunteer Rescue Squad.
Ethical Statement
This research was reviewed and approved by the Institutional Review Board for the Social and Behavioral Sciences (IRB-SBS) at the University of Virginia, which established protocols for participant recruitment, informed consent, experimental procedures, and data management. All participants were fully informed of the study’s purpose, procedures, potential risks and benefits, and future data usage through consent and materials release forms. Only data from participants who provided explicit consent for de-identified publication are included in the dataset. In accordance with the IRB protocol, all identifying information was removed prior to release. Each participant was assigned a unique identifier, faces were blurred semi-automatically in egocentric video, and sensitive textual or audio identifiers such as names, license plates, and ID cards were manually obscured. All data underwent manual verification to ensure privacy preservation and compliance with ethical standards.
EgoEMS was developed using simulated emergency scenarios with trained responders under IRB oversight, enabling high-fidelity modeling of real-world conditions while maintaining participant safety and privacy. However, the extension of such systems to real-world EMS introduces substantial ethical and privacy challenges. Obtaining informed consent during emergencies is often infeasible when patients are unconscious or in critical conditions, and incidentally captured bystanders or minors may not have provided consent. Real scenes also encompass sensitive situations such as domestic violence, substance use, or mental health crises, which require enhanced safeguards, institutional oversight, and alignment with frameworks like HIPAA. Furthermore, ensuring robust de-identification and responsible data governance in unconstrained environments remains an open technical and ethical challenge.
Beyond data collection, the deployment of ICAs in EMS contexts raises broader societal considerations regarding surveillance, bias, and trust. Models trained on simulated data may underperform in the field due to shifts in environmental and behavioral distributions, necessitating rigorous validation, transparency, and accountability measures before clinical use. Our work therefore positions EgoEMS as a privacy conscious proof-of-concept to advance ICAs responsibly, by offering an open-source data collection framework, EMS-specific ontology, annotation tools, and de-identification pipeline to support gradual, ethically governed expansion to real-world settings.
We advocate for future efforts to establish consent and governance frameworks in collaboration with IRBs, EMS agencies, and legal experts; to develop policy-level standards inspired by body-worn camera protocols; and to implement on-device, HIPAA-compliant data processing with built-in de-identification. At the algorithmic level, we emphasize the importance of multimodal learning methods that prioritize privacy-preserving modalities (e.g., IMU or depth sensors) when visual data is restricted. We also propose using transfer learning methods and exploiting multimodal redundancy to reduce the sim-to-real gap and improve models robustness and generalizability. Together, these principles aim to ensure that cognitive assistance technologies for emergency response evolve safely, fairly, and in service of public trust and societal benefit.
@@ A.1 NEMSIS
check_pulse Assess for carotid or radial pulse.
check_breathing Look, listen, and feel for breathing.
chest_compressions Perform compressions at correct rate/depth.
assess_patient Reassess patient condition.
Defibrillation request_aed Request AED early in resuscitation.
request_assistance Call for bystander or team assistance.
turn_on_aed Power on AED and prepare for use.
attach_defib_pads Attach pads to bare chest correctly.
clear_for_analysis Ensure no contact during rhythm analysis.
clear_for_shock Confirm everyone is clear before shock.
administer_shock_aed Deliver shock if advised by AED.
Ventilation open_airway Use airway technique (e.g., head-tilt).
place_bvm Position and seal BVM mask.
ventilate_patient Deliver ventilations appropriately.
suction_airway Use suction to clear airway.
inset_NPA Insert NPA to maintain airway.
Intervention Keystep Description
@@ A.2 Keysteps
Following selecting the protocols and interventions, we consulted the National Registry of Emergency Medical Technicians (NREMT) to review the keysteps and guidelines necessary for accurately executing each intervention. To ensure the robustness of our taxonomy, we collaborate closely with EMS experts, who validate the defined keysteps for each procedure. Specifically, we reference the Emergency Medical Responder Psychomotor Examination criteria nremt_psychomotor_exams to inform the development of this keystep taxonomy, aligning it with established standards for practical skill assessment. Tables 4,5,6 present the taxonomy of EMS interventions and keysteps defined for our dataset.
@@ A.3 Scenarios
This structure cascades into specific interventions (e.g., Vital Sign Assessment, 12-lead ECG, Transport), each decomposed into granular keysteps (e.g., explain procedure, place V1 lead, interpret ECG). To support model generalization and realism, we designed diverse scenarios featuring patients with varying comorbidities and sociocultural backgrounds. Scenario briefs were provided to participants before each simulation to standardize objectives while allowing natural variations and reactions in responder-patient interactions. For example, in the context of cardiac arrest emergency, we created multiple scenarios (see Figure 8) featuring patients with varied backgrounds and comorbidities, thereby enhancing the diversity and ecological validity of the dataset.
@@ A.4.1 Keystep Annotations
Figure 10: Video annotation tool from Oxford dutta2019vgg.
@@ D.1.1 Fully Supervised Setting
Learning Type Model (Architecture) Modality (Feature Extraction) Classification Metric
Precision Recall F1-score Top-1 Acc. Top-5 Acc.
Supervised MTRSAP (Transformer) weerasinghe2024multimodal Ego Video (Resnet50) 0.55 0.58 0.55 62.27 77.01
Ego Video (I3D) 0.56 0.57 0.56 60.41 72.50
Ego Video, IMU (Resnet50, Norm.) 0.55 0.58 0.56 62.19 76.14
IMU (Norm.) 0.42 0.45 0.42 48.34 67.57
Ego Video, Audio (Resnet50, Mel-Spec) 0.19 0.23 0.20 25.67 44.38
Ego Video, Audio (Resnet50, Wav2Vec2) 0.34 0.39 0.35 44.71 66.16
Audio (Wav2Vec2) 0.26 0.27 0.26 33.25 52.14
Few-Shot MM-CDFSL (Transformer) Hatano2024MMCDFSL Ego Video (Multimodal Distilled) $\dagger$ - - - 67.70 -
Zero-Shot VideoLlama-3.3 (VLM) zhang2025videollama Ego Video (End-to-End) 0.16 0.17 0.16 18.51 -
Qwen-2.5 (VLM) bai2025qwen2 Ego Video (End-to-End) 0.33 0.35 0.34 38.30 -
GPT-4o (LLM) achiam2023gpt Audio (End-to-End) 0.17 0.19 0.18 23.89 -
@@ D.1.2 Multimodal Few-Shot Setting
Knowledge transfer and multimodal distillation show promising performance in low-data regimes. To assess cross-domain transfer, we evaluate cross-domain, multimodal few-shot learning with MM-CDFSL Hatano2024MMCDFSL. Following the authors, we use the distilled RGB student encoder on a large scale general purpose egocentric dataset (Ego4D grauman2022ego4d) (teachers: optical flow and hand pose) and perform episodic 5-way, 5-shot evaluation on our dataset. For each episode, a linear classifier (logistic regression) is fit on the support set and evaluated on the query set using RGB only; we also apply ensemble masked inference as in the original work. We report the mean top-1 accuracy across 600 episodes.